If z=3+2i evaluate z^4-z^3+z^2-z+1
Answer by Alan3354(69443)

If z=3+2i evaluate z^4-z^3+z^2-z+1
----------
Step 1, find z^2
z^2 = (3+2i)*(3+2i) = 9 + 12i + 4i^2 = 9 + 12i - 4 = 5 + 12i
----
Multiply that by z to find z^3
----
Square it to find z^4
---
Add them.